OpenStudy (anonymous):

what is 19= h/3 - 8? and explain how you got it

OpenStudy (anonymous):

19=h/3 - 8\[19+8=(h/3)-8+8\] adding 8 to both the sides \[27=h/3\] multiplying both sides by 3 \[27*3=(h/3)*3\] equation simplifies and value of h is obtained \[81=h\]